OpenStudy (gabbyalicorn):

----- What is ^3√-125

OpenStudy (gabbyalicorn):

@sammixboo please help me

OpenStudy (lapinblanc):

Are you dealing with imaginary numbers right now in your work?

OpenStudy (gabbyalicorn):

I'm not sure what that means, but im just dealing with irrational numbers.

OpenStudy (lapinblanc):

And did you mean to have the carrot in that questions. (the ^ symbol)

OpenStudy (lapinblanc):

Well in order to take the root of a negative number, you use imaginary numbers, or i.

OpenStudy (lapinblanc):

Ah. the three is the root for the square.

OpenStudy (gabbyalicorn):

let me write it out

OpenStudy (lapinblanc):

well, 125^(1/3) is 5*5*5

OpenStudy (lapinblanc):

And because it is negative, you just put an i after the number, 5i

OpenStudy (lapinblanc):

because the third root of 125 is 5.

OpenStudy (lapinblanc):

Answer: 5i

OpenStudy (gabbyalicorn):

i dont think i could put i in my fill in the blank

OpenStudy (lapinblanc):

Does your question provide a link to look at the chapter your are doing work in? That will tell me if you are doing complex or just irrational. The only answer to this question has an i because of that negative number under the roof.

OpenStudy (gabbyalicorn):

just irrational

OpenStudy (agent0smith):

@lapinblanc only even roots are imaginary. If the exponent is odd, then the root is real.

OpenStudy (agent0smith):

Since (-5)(-5)(-5) = -125 So the cube root of -125 is -5.

OpenStudy (lapinblanc):

Silly me, because -5*-5*-5.

OpenStudy (gabbyalicorn):

\(º-º)/

OpenStudy (gabbyalicorn):

Thank you @agent0smith

OpenStudy (lapinblanc):

When you take the root of something, your root tells you how many times it has to be repeated to remove it. for example \[\sqrt[2]{4}\] Is the same thing as saying \[\sqrt[2]{2*2}\] So the answer is 2. In our case we have \[\sqrt[3]{-125}\] which is the same as \[\sqrt[3]{-5*-5*-5}\] so our answer is -5

OpenStudy (zzr0ck3r):

\(\sqrt[\text{odd}]{-a}=-\sqrt[\text{odd}]{a}\)

OpenStudy (gabbyalicorn):

Oh, i think i get it. :D So we just needed to find the number that would go in 3 times. I see. :)

OpenStudy (lapinblanc):

Precisely.

OpenStudy (gabbyalicorn):

Wow, simpler than i had thought

OpenStudy (gabbyalicorn):

Thank you so much for you're help. :) It's really appreciated

OpenStudy (lapinblanc):

^__^